A house with an original value of $180,000. increased in vlue to $220,000. in 5 years. What is the ratio of the increase in price to the original price.

$220,000/$180,000 = 11/9

Answer by stanbon(75887)

A house with an original value of $180,000. increased in vlue to $220,000. in 5 years. What is the ratio of the increase in price to the original price.
-------
Increase in price = (220,000-180,000) = 40,000
---
Ratio = 40,000/180,000 = 40/180 = 2/9 = 2:9
================================================
Cheers,
Stan H.